OpenStudy (anonymous):

find dy/dx for (sqrt(x)-1)/(sqrt(x) +1)

OpenStudy (anonymous):

if possible could you assist me @ganeshie8 i would appreciate it

OpenStudy (anonymous):

could you explain how you got that answer for the derivative of [(sqrt{x}-1)/(sqrt{x}+1)

OpenStudy (anonymous):

look the power rule says: func) x^n = Deri) nx^n−1

OpenStudy (anonymous):

nx^n-1

OpenStudy (anonymous):

i understand that but could you explain it in steps as i am confused where to apply the power rule for this equation

OpenStudy (anonymous):

So \[\sqrt x\] can be written as \[\huge x^{1\over2} \]

OpenStudy (anonymous):

yes i understand that

OpenStudy (anonymous):

So you differentiate it, and show up!

OpenStudy (anonymous):

thank you for your time and effort @Rohangrr i understand now

OpenStudy (anonymous):

Thanks, hope it helped. Any more help needed when I'm offline just mail @ rcdrohan@gmail.com

OpenStudy (anonymous):

Arigato ozaimas ~

OpenStudy (unklerhaukus):

For a function of the form \[y(x)=\frac{f(x)}{g(x)}\]the quotient rule says\[y'(x)=\left(\frac{f}{g}\right)'=\frac{f'g-fg'}{g^2}\] \[y(x) = \frac{\sqrt x-1}{\sqrt x +1}\]\[\frac{\mathrm dy}{\mathrm dx} = \frac{\frac{\mathrm d}{\mathrm dx}(\sqrt x-1)\cdot(\sqrt x+1)-(\sqrt x-1)\cdot \frac{\mathrm d}{\mathrm dx}(\sqrt x+1)}{(\sqrt x+1)^2}\] It might me easiest to calculate \[f'=\frac{\mathrm d}{\mathrm dx}(\sqrt x-1)\] and\[g'=\frac{\mathrm d}{\mathrm dx}(\sqrt x+1)\]first

OpenStudy (unklerhaukus):

\[f'=\frac{\mathrm d}{\mathrm dx}(\sqrt x-1)\\\quad=\frac{\mathrm d}{\mathrm dx}(x^{1/2}-x^0)\\\quad=\tfrac12x^{1/2-1}-0\, x^{0-1}\\\quad=\tfrac12x^{-1/2}\\\quad=\frac1{2\sqrt x}\]

OpenStudy (anonymous):

thank you @UnkleRhaukus for your help